OpenStudy (anonymous):

Pre calc: simplify tan^2x cos^x + cot^2x sin^2x

OpenStudy (anonymous):

\[\tan ^2x \cos^2x+\cot^2x \sin^2x\]

OpenStudy (welshfella):

try converting tan and cot to sines and cosines

OpenStudy (welshfella):

- and the answer will be a number!

OpenStudy (anonymous):

Before I write it out, i have a question. I know tan= sin/cos so i figured tan^2 would equal (sin/cos)^2. but Symolab is telling me tan^2= (cos/sin)^2 . is that right?

OpenStudy (welshfella):

tan^2 x = sin^2 x / cos^2 x

OpenStudy (welshfella):

- which is another way of writing (sin x / cos x)^2

OpenStudy (anonymous):

\[\frac{ \sin^2x }{ \cos^2x } \cos^2x + \frac{ \cos^2x }{ \sin^2x } \sin^2x\]

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Do I distribute the cos^2x and the sin^2x to their respective sides now?

OpenStudy (mikwind):

use the same identity from your other problem :D

OpenStudy (welshfella):

you can quickly simplify this . Look at the occurences of sin^2x and cos^2 x

OpenStudy (welshfella):

compare